# Quantum-Enhanced Processing with Tensor-Network Frontends for Privacy-Aware Federated Medical Diagnosis

### 03 Apr 2026 v2

@@ I Introduction
MPC strengthens privacy protection but introduces a systems-level trade-off, because protecting richer intermediate representations generally increases communication and execution cost. This trade-off becomes more restrictive when a quantum post-processing module is introduced, since the quantum branch must operate on whatever protected latent representation remains practical after secure aggregation. Meanwhile, prior work on distributed and hybrid quantum learning has studied distributed quantum models [26, 31, 2, 12], tensor-network/quantum hybrid architectures [6, 21, 34, 7, 20, 32], and privacy-preserving quantum federated protocols [19]. However, these lines of work have typically focused on protocol design or model design separately, and they do not sufficiently characterize how post-aggregation quantum refinement behaves when jointly constrained by secure aggregation and representation compression.
@@ II-A Client-Side Tensor-Network Embedding
For the TTN and MERA branches, the input is first reorganized into non-overlapping image patches. In particular, the $28\times 28$ image is partitioned into 1616 patches of size $7\times 7$, and each patch is processed by a shared patch-wise stem network:
@@ II-A3 MERA
II-B Security Model and MPC Preliminaries
II-B1 System Model and Threat Model
We consider $n$ clients $\{C_{i}\}_{i=1}^{n}$ and an outsourced aggregation server implemented by three non-colluding computation nodes $S_{0},S_{1},S_{2}$. Each client holds a private dataset $\mathcal{D}_{i}$ and computes a local representation $f_{i}\in\mathbb{R}^{d}$; for the MPC benchmark, a client-specific scalar coefficient $w_{i}\in\mathbb{R}$ is additionally introduced to model protected weighted aggregation and normalization. We consider both passive and active security, corresponding to Scenarios 1–3 and Scenarios 4–6, respectively.
II-B2 Cost Model
II-C Secure Aggregation via MPC
@@ II-D1 Quantum Encoding and Observable Readout
Figure 2: Parameterized quantum circuit used in the QEP. Each layer applies single-qubit $R_{y}$ and $R_{z}$ rotations followed by nearest-neighbor CNOT entangling gates. For visualization, terminal measurement gates are shown explicitly. In the actual implementation, however, the circuit output is summarized through expectation values of selected Pauli observables rather than bitstring sampling.
@@
where $N_{q}$ is the number of qubits and $\mathrm{Enc}(\cdot)$ denotes a shallow multilayer perceptron with normalization and nonlinear activation. For each layer $l$ and qubit $q$, rotation angles are defined as
@@
with global scale $s$ and circuit-side angle offsets $\delta^{(l,q)}_{y},\delta^{(l,q)}_{z}$.
@@
For a circuit with $N_{q}$ qubits and depth $L$, this construction uses $2LN_{q}$ single-qubit rotation gates and $L(N_{q}-1)$ nearest-neighbor CNOT gates. The final quantum feature dimension is then determined by the selected observable set used for expectation-value evaluation. Fig. 2 illustrates the circuit structure used in the QEP.
The resulting state is summarized through expectation values of Pauli observables. For each observable $\hat{O}_{m}$, the corresponding quantum statistic is
@@
which yields a raw quantum feature vector
@@
where $d_{q}$ is determined by the number and type of observables. This yields observable-based quantum statistics from a compact latent representation without requiring a large quantum register.
@@ III Experimental Setup
We evaluated the proposed framework on PneumoniaMNIST, a binary medical image classification benchmark from the MedMNIST collection [36]. The task is to discriminate Normal from Pneumonia using resized $28\times 28$ grayscale chest X-ray images normalized to $[0,1]$. For the MPS branch, each image was flattened into a 784-dimensional vector. For the TTN and MERA branches, the same image was additionally reorganized into 1616 non-overlapping patches of size $7\times 7$, followed by a shared patch-wise stem transformation. The standard training, validation, and test splits provided by the dataset interface were used throughout.
@@ IV-B Experiment 2: MPC Mode Benchmark
We vary the number of clients from $n=1$ to 3030. For each value of $n$, we consider seven scenarios. Scenario 0 is an insecure baseline in which a central server $S$ performs the aggregation and transformation directly. Scenarios 1–3 correspond to passive-security MPC settings with progressively richer protected functionality: Scenario 1 protects only weighted aggregation, Scenario 2 additionally protects normalization, and Scenario 3 further protects the post-aggregation transformation stage. Scenarios 4–6 are the corresponding active-security counterparts. Table I summarizes these settings.
@@ VI Discussion
The results of Experiment 1 suggest that the QEP should be understood not as a universal accuracy booster, but as a quantum feature-refinement mechanism whose effectiveness depends on how it is paired with the tensor-network frontend. Across MPS, TTN, and MERA, the QEP does not yield a uniform increase in Accuracy or F1-score. Instead, its contribution appears through architecture-dependent changes in class-wise operating characteristics and in the internal fusion dynamics of the model. The main implication is therefore not the superiority of any single component in isolation, but the importance of jointly designing tensor-network structure and post-aggregation quantum refinement.
Within this broader picture, TTN exhibits the most favorable overall profile in the present setting. It achieves strong predictive performance, including high Accuracy and Pneumonia F1, while maintaining high sensitivity to pneumonia-positive cases. More importantly, however, the results indicate that the contribution of the QEP is conditioned by the latent representation supplied by the frontend. MPS, TTN, and MERA do not merely compress the input differently; they expose different latent organizations to the QEP, which in turn leads to different quantum operating regimes.
The internal diagnostics support this interpretation. MPS exhibits the largest quantum-branch dispersion, whereas TTN remains in a lower and more controlled $q$-std regime. This suggests that larger internal quantum variability does not automatically translate into stronger downstream performance. In the present experiment, the most favorable learning behavior emerges not from the largest quantum-branch spread, but from a more balanced interaction between quantum refinement and the frontend-induced latent structure. From a quantum-machine-learning perspective, this indicates that the effect of post-aggregation quantum enhancement is jointly determined by the quantum map and the representational topology that feeds it.
The qubit-scaling and noise analyses provide an additional practical perspective on the TTN-based QEP. For the present 64-dimensional latent input, lower-qubit configurations, particularly $N_{q}=4$ and $N_{q}=6$, appear to provide insufficient feature capacity, whereas the behavior becomes more stable from $N_{q}=8$ onward. In addition, all noisy conditions at $N_{q}=8$ show some degree of degradation relative to the noiseless case, although the extent of degradation depends on the noise model. These observations do not establish hardware scalability, but they indicate that, in the TTN setting, the hybrid pipeline retains meaningful operating behavior once the quantum register size is reasonably matched to the latent dimension.
@@
At the same time, the present results should not be read as establishing TTN+QEP as a universally optimal combination. Rather, they indicate that the utility of post-aggregation quantum refinement is contingent on the latent structure delivered by the frontend and on how that structure matches the effective capacity of the quantum feature map. In this sense, the main lesson is not the dominance of a single architecture, but the importance of co-design across compression, aggregation, and quantum refinement.
@@
Experiment 2 complements this interpretation from the systems side. Because the communication cost is dominated by the protected representation dimension, tensor-network compression directly reduces the communication overhead incurred by MPC on that representation. Thus, the framework not only shapes the latent interface for the QEP, but also improves the practical outlook for privacy-preserving deployment by lowering the effective cost of secure aggregation.
@@
A related direction for quantum feature extraction in this framework is to replace or augment the present observable-based QEP with a quantum reservoir readout acting on the aggregated latent input. Quantum reservoir computing is attractive in this context because it exploits the natural nonlinear dynamics of many-body quantum systems while avoiding gradient-based variational optimization of the quantum circuit itself. This is relevant from a practical viewpoint, as variational quantum models can suffer from barren plateau phenomena, whereas quantum kernel methods can also face exponential concentration of kernel values and associated measurement inefficiency under certain conditions [18, 33]. By contrast, reservoir-style quantum processing uses the dynamics of a fixed quantum system as a feature generator and trains only the classical readout, which provides a complementary design point for small- to intermediate-scale hybrid learning. Recent studies have experimentally demonstrated repeated-measurement quantum reservoir computing on superconducting devices and have also reported large-scale analog quantum reservoir learning at the 100-qubit scale, supporting the practical relevance of this direction [37].
@@
The longer-term significance of the architecture lies in future qubit regimes where faithful classical simulation is no longer practical, in which case post-aggregation quantum processing on tensor-network-compressed latent features becomes a more meaningful systems option than direct quantum processing of the original high-dimensional input. This interpretation is also consistent with the qubit-scaling result. In the present setting, improved stability emerges only once the quantum feature capacity becomes reasonably matched to the latent representation size, suggesting that frontend structure and quantum resource scale should be considered jointly rather than independently.
@@
At the same time, the present study has several limitations. First, the quantum branch is evaluated through Qiskit Aer statevector simulation, so the results should be interpreted as evidence for hybrid representational behavior rather than as a demonstration of hardware-level feasibility. Moreover, the present study does not establish that the quantities computed by the QEP are classically intractable or impractical to obtain classically. Rather, the QEP is used here as a structured small-qubit feature map, and the focus is on its representational interaction with tensor-network-compressed latent features rather than on computational advantage. Second, the current comparison does not yet isolate which portion of the observed QEP effect is uniquely attributable to the quantum observable map itself, as opposed to the broader effect of introducing an additional nonlinear transformation pathway. Third, the MPC benchmark is based on symbolic communication-cost modeling and therefore does not replace full protocol-level implementation and empirical measurement.
@@ VII Conclusion
The additional qubit-scaling and noise analyses further clarify the practical behavior of the QEP. For the present 64-dimensional latent input, performance becomes more stable from 88 qubits onward, while noisy conditions at 88 qubits degrade performance relative to the noiseless case in a noise-type-dependent manner. These results do not establish hardware-level benefit, but they help characterize the operating range and robustness of the hybrid quantum branch within the proposed pipeline.
@@ Acknowledgments
The authors acknowledge helpful conversations with Mark Medum Bundgaad, Miyoji Kakinuki, Akiko Kamigori, Kazufumi Okazaki, Takuya Hasegawa, Tomah Sogabe, Naoki Yamamoto, Eriko Kaminishi, Toshiki Yasuda. Hiroshi Yamauchi acknowledges the support of Yoshi-aki Shimada, Yosuke Komiyama and Ryuji Wakikawa. This work was partly supported by NEDO Challenge Quantum Computing “Solve Social Issues!”.

### 13 Apr 2026 v3

@@ I Introduction
MPC strengthens privacy protection but introduces a systems-level trade-off, because protecting richer intermediate representations generally increases communication and execution cost. This trade-off becomes more restrictive when a quantum post-processing module is introduced, since the quantum branch must operate on whatever protected latent representation remains practical after secure aggregation. Meanwhile, prior work on distributed and hybrid quantum learning has studied distributed quantum models [29, 35, 3, 15], tensor-network/quantum hybrid architectures [8, 24, 38, 9, 23, 36], privacy-preserving quantum federated protocols [22], and quantum-secure secure-computation schemes for deep learning inference that bound leakage of both client data and model weights [34]. However, these lines of work have typically focused on protocol design, secure inference, or model design separately, and they do not sufficiently characterize how post-aggregation quantum refinement behaves when jointly constrained by secure aggregation and representation compression. This distinction is important for our setting: rather than replacing the federated pipeline with a fully quantum-secure inference engine, we study how representation compression and small-qubit post-aggregation quantum processing can be integrated with an MPC-based secure-aggregation model under practical deployment constraints.
@@ II-A Client-Side Tensor-Network Embedding
For the TTN and MERA branches, the input is first reorganized into non-overlapping image patches. In particular, the $28\times 28$ image is partitioned into $16$ patches of size $7\times 7$, and each patch is processed by a shared patch-wise stem network:
@@ II-A3 MERA
II-B Secure Aggregation via MPC
@@
II-C Security Model and MPC Preliminaries
II-C1 System Model and Threat Model
We consider $n$ clients $\{C_{i}\}_{i=1}^{n}$ and an outsourced aggregation server implemented by three non-colluding computation nodes $S_{0},S_{1},S_{2}$. Each client holds a private dataset $\mathcal{D}_{i}$ and computes a local representation $f_{i}\in\mathbb{R}^{d}$; for the MPC benchmark, a client-specific scalar coefficient $w_{i}\in\mathbb{R}$ is additionally introduced to model protected weighted aggregation and normalization. We consider both passive and active security.
II-C2 MPC Building Blocks
The scenarios considered later are modeled using the MPC framework of [1], together with standard techniques for secure fixed-point computation over real values [11]. This subsection briefly summarizes the relevant building blocks.
Secret sharing
A value $v\in\mathbb{Z}_{2^{k}}$ is shared as follows. Sample $v_{0},v_{1}$ uniformly at random from $\mathbb{Z}_{2^{k}}$ and define $v_{2}=v-v_{0}-v_{1}$. Let $v]_{i}=(v_{i},v_{i+1\,\mathrm{mod}\,3})$ denote the share held by computation node $S_{i}$ for $i\in\{0,1,2\}$. In what follows, subscripts are implicitly taken modulo 3. This sharing scheme has two key properties: (1) any single share pair reveals no information about $v$; and (2) any two nodes can reconstruct $v$ by combining all three components and computing $v=v_{0}+v_{1}+v_{2}$ This scheme is known as replicated secret sharing.
Given secret sharings $v]_{i}=(v_{i},v_{i+1})$ and $w]_{i}=(w_{i},w_{i+1})$, the pair $(v_{i}+w_{i},\;v_{i+1}+w_{i+1})$ is a valid secret sharing of $v+w$. By similar reasoning, subtraction and multiplication by public constants can also be performed locally.
To multiply $[v]$ and $[w]$, each node $S_{i}$ first computes
$z_{i}=v_{i}w_{i}+v_{i+1}w_{i}+v_{i}w_{i+1}$ (23)
and then sends $z_{i}$ to $S_{i-1}$. Each node finally defines $z]_{i}=(z_{i},z_{i+1})$, which is a valid secret sharing of vwvw. [^1] This description omits additional correction terms required for full security. These terms do not affect the communication complexity and are therefore omitted for clarity. Thus, one secure multiplication requires the exchange of three $k$-bit elements.
Active security can be achieved as in [1] by effectively running the protocol twice in parallel: once on the real inputs, and once on a suitably randomized encoding of those inputs. A final consistency check is then performed to detect deviations.
II-C3 Fixed-Point and Truncation Semantics
To support computation over real numbers, values are encoded in fixed-point representation. Concretely, a real value $v\in\mathbb{R}$ is represented as
$\tilde{v}=\lfloor v\cdot 2^{F}\rceil\in\mathbb{Z}_{2^{k}},$ (24)
where $F$ is the fractional precision parameter. Under this encoding, multiplication yields a value scaled by $2^{2F}$, so a truncation step is required to restore the original scale. This truncation incurs additional communication overhead: specifically, six $k$-bit elements must be exchanged [11].
II-C4 Division of Secret-Shared Values
The normalization in (22) requires division of two secret-shared quantities, namely $\mathrm{WF}$ and $\mathrm{W}+\varepsilon$. This can be implemented using Goldschmidt division as described in [7]. At a high level, the method proceeds in two stages. First, an initial approximation to $1/(\mathrm{W}+\varepsilon)$ is computed. Second, an iterative refinement is performed to approximate $\mathrm{WF}/(\mathrm{W}+\varepsilon)$. The first stage requires a truncation step and a secure computation of the effective bit width of $\mathrm{W}+\varepsilon$. Determining the bit width of a secret-shared value requires bit decomposition, which can be implemented using approximately $k$ secure multiplications. The second stage requires two fixed-point multiplications per iteration. The required number of iterations depends on the ring size.
II-C5 Cost Model
@@ II-D1 Quantum Encoding and Observable Readout
Figure 2: Parameterized quantum circuit used in the QEP. Each layer applies single-qubit $R_{y}$ and $R_{z}$ rotations followed by nearest-neighbor CNOT entangling gates. For visualization, the circuit diagram is shown in a 4-qubit instance, and terminal measurement gates are drawn explicitly. In the actual implementation, however, the circuit output is summarized through expectation values of selected Pauli observables rather than bitstring sampling. In the experiments, the same circuit template was evaluated with different qubit counts, up to $N_{q}=16$.
@@
where $N_{q}$ is the number of qubits and $\mathrm{Enc}(\cdot)$ denotes a shallow multilayer perceptron with normalization and nonlinear activation. For each layer $l$ and qubit $q$, the rotation angles are defined as
@@
where $s$ is a global scaling factor and $\delta^{(l,q)}_{y},\delta^{(l,q)}_{z}$ are circuit-side offsets.
@@
For a circuit with $N_{q}$ qubits and depth $L$, this construction requires $2LN_{q}$ single-qubit rotation gates and $L(N_{q}-1)$ nearest-neighbor CNOT gates. Fig. 2 illustrates the circuit structure used in the QEP.
This leads to an overall gate complexity of $\mathcal{O}(LN_{q})$, with nearest-neighbor entangling gates. In contrast, when both one-body and two-body observables are included, the effective quantum feature dimension scales as $d_{q}=\mathcal{O}(N_{q}^{2})$. Thus, the circuit realizes a quadratic feature expansion while maintaining linear gate complexity.
The output quantum state is summarized through expectation values of Pauli observables. For each observable $\hat{O}_{m}$, the corresponding quantum statistic is defined as
@@
yielding a feature vector
@@
where $d_{q}$ is determined by the chosen observable set. This observable-based readout enables extraction of nonlinear quantum features from a compact latent representation, as only expectation values of selected observables are evaluated. This construction should be interpreted as an observable-based quantum feature map rather than a variational, kernel-based, or tomography-based approach.
@@ III Experimental Setup
We evaluated the proposed framework on PneumoniaMNIST, a binary medical image classification benchmark from the MedMNIST collection [40]. The task is to discriminate Normal from Pneumonia using resized $28\times 28$ grayscale chest X-ray images normalized to $[0,1]$. For the MPS branch, each image was flattened into a 784-dimensional vector. For the TTN and MERA branches, the same image was additionally reorganized into $16$ non-overlapping patches of size $7\times 7$, followed by a shared patch-wise stem transformation. The standard training, validation, and test splits provided by the dataset interface were used throughout.
@@ IV-B Experiment 2: MPC Mode Benchmark
We vary the number of clients from $n=1$ to $30$. For each value of $n$, we consider seven scenarios. Scenario 0 is an insecure baseline in which a central server $S$ performs the aggregation and transformation directly. Scenarios 1–3 correspond to passive-security MPC settings with progressively richer protected functionality: Scenario 1 protects only weighted aggregation, Scenario 2 additionally protects normalization, and Scenario 3 further protects the post-aggregation transformation stage. Scenarios 4–6 are the corresponding active-security counterparts. Table I summarizes these settings.
@@ VI Discussion
The results of Experiment 1 suggest that the QEP should be understood not as a universal accuracy booster, but as a quantum feature-refinement mechanism whose effectiveness depends on how it is paired with the tensor-network frontend. Across MPS, TTN, and MERA, the QEP does not yield a uniform increase in Accuracy or F1-score; instead, its contribution appears through architecture-dependent changes in class-wise operating characteristics and internal fusion dynamics. The main implication is therefore not the superiority of any single component in isolation, but the importance of jointly designing tensor-network structure and post-aggregation quantum refinement.
Within this broader picture, TTN exhibits the most favorable overall profile, achieving strong predictive performance, including high Accuracy and Pneumonia F1, while maintaining high sensitivity to pneumonia-positive cases. More importantly, the results indicate that the contribution of the QEP is conditioned by the latent representation supplied by the frontend. MPS, TTN, and MERA do not merely compress the input differently; they expose different latent organizations to the QEP, leading to distinct quantum operating regimes.
The internal diagnostics support this interpretation. MPS exhibits the largest quantum-branch dispersion, whereas TTN remains in a lower and more controlled $q$-std regime, suggesting that larger internal quantum variability does not automatically translate into stronger downstream performance. In the present experiment, the most favorable behavior emerges not from the largest quantum-branch spread, but from a more balanced interaction between quantum refinement and the frontend-induced latent structure. From a quantum-machine-learning perspective, this indicates that the effect of post-aggregation quantum enhancement is jointly determined by the quantum map and the representational topology that feeds it.
The qubit-scaling and noise analyses provide an additional practical perspective on the TTN-based QEP. For the present 64-dimensional latent input, lower-qubit configurations, particularly $N_{q}=4$ and $N_{q}=6$, appear to provide insufficient feature capacity, whereas behavior becomes more stable from $N_{q}=8$ onward. In addition, all noisy conditions at $N_{q}=8$ show some degradation relative to the noiseless setting, although the extent depends on the noise model. These observations do not establish hardware scalability, but indicate that the hybrid pipeline retains meaningful behavior once the quantum register size is reasonably matched to the latent dimension.
@@
At the same time, the present results should not be read as establishing TTN+QEP as a universally optimal combination. Rather, they indicate that the utility of post-aggregation quantum refinement depends on the latent structure delivered by the frontend and its alignment with the effective capacity of the quantum feature map. In this sense, the main lesson is not the dominance of a single architecture, but the importance of co-design across compression, aggregation, and quantum refinement.
@@
Experiment 2 complements this interpretation from the systems side. Because communication cost is dominated by the protected representation dimension, tensor-network compression directly reduces the communication overhead incurred by MPC on that representation, improving the practicality of privacy-preserving deployment.
@@
A related direction for quantum feature extraction in this framework is to replace or augment the present observable-based QEP with a quantum reservoir readout acting on the aggregated latent input. Quantum reservoir computing is attractive in this context because it exploits the natural nonlinear dynamics of many-body quantum systems while avoiding gradient-based variational optimization of the quantum circuit itself. This is relevant from a practical viewpoint, as variational quantum models can suffer from barren plateau phenomena, whereas quantum kernel methods can also face exponential concentration of kernel values and associated measurement inefficiency under certain conditions [21, 37]. By contrast, reservoir-style quantum processing uses the dynamics of a fixed quantum system as a feature generator and trains only the classical readout, providing a complementary design point for small- to intermediate-scale hybrid learning. Recent studies have experimentally demonstrated repeated-measurement quantum reservoir computing on superconducting devices and large-scale analog quantum reservoir learning at the 100-qubit scale, supporting the practical relevance of this direction [41].
@@
The longer-term significance of the architecture lies in future qubit regimes where faithful classical simulation is no longer practical, in which case post-aggregation quantum processing on tensor-network-compressed latent features becomes more meaningful than direct processing of the original high-dimensional input. This interpretation is consistent with the qubit-scaling result, suggesting that frontend structure and quantum resource scale should be considered jointly.
@@
At the same time, the present study has several limitations. First, the quantum branch is evaluated through Qiskit Aer statevector simulation, so the results should be interpreted as evidence for hybrid representational behavior rather than as a demonstration of hardware-level feasibility. Second, the present study does not establish that the quantities computed by the QEP are classically intractable or impractical to obtain classically; rather, the QEP is used as a structured small-qubit feature map. Third, the current comparison does not yet isolate which portion of the observed QEP effect is uniquely attributable to the quantum observable map itself, and the MPC benchmark remains a symbolic communication-cost model rather than a full protocol-level implementation.
@@ VII Conclusion
The additional qubit-scaling and noise analyses further clarify the practical behavior of the QEP. For the present 64-dimensional latent input, performance becomes more stable from $8$ qubits onward, while noisy conditions at $8$ qubits degrade performance relative to the noiseless case in a noise-type-dependent manner. These results do not establish hardware-level benefit, but they help characterize the operating range and robustness of the hybrid quantum branch within the proposed pipeline.
@@ Acknowledgments
The authors acknowledge helpful conversations with Mark Medum Bundgaad, Miyoji Kakinuki, Akiko Kamigori, Kazufumi Okazaki, Takuya Hasegawa, Tomah Sogabe, Naoki Yamamoto, Eriko Kaminishi, Toshiki Yasuda. Hiroshi Yamauchi acknowledges the support of Yoshi-aki Shimada, Yosuke Komiyama and Ryuji Wakikawa. This work was partly supported by NEDO Challenge Quantum Computing “Solve Social Issues!”. Portions of the manuscript were initially prepared in Japanese and subsequently refined in English using AI-based machine translation tools. All content was subsequently reviewed and approved by all authors.